Lawn Spray Inc. develops and produces spraying equipment for lawn maintenance and industrial uses. On January...

  1. Lawn Spray Inc. develops and produces spraying equipment for lawn maintenance and industrial uses. On January 31 of the current year, Lawn Spray Inc. reacquired 8,800 shares of its The stock outstanding when a corporation has issued only one class of stock.common stock at $29 per share. On June 14, 5,500 of the reacquired shares were sold at $31 per share, and on November 23, 2,100 of the reacquired shares were sold at $33.

    a. Journalize the transactions of January 31, June 14, and November 23. For a compound transaction, if an amount box does not require an entry, leave it blank.

    Jan. 31
    • Cash
    • Common Stock
    • Paid-In Capital from Sale of Treasury Stock
    • Preferred Stock
    • Treasury Stock
    • Cash
    • Common Stock
    • Paid-In Capital from Sale of Treasury Stock
    • Retained Earnings
    • Treasury Stock
    June. 14
    • Cash
    • Common Stock
    • Paid-In Capital from Sale of Treasury Stock
    • Retained Earnings
    • Treasury Stock
    • Cash
    • Common Stock
    • Preferred Stock
    • Retained Earnings
    • Treasury Stock
    • Cash
    • Common Stock
    • Paid-In Capital from Sale of Treasury Stock
    • Preferred Stock
    • Retained Earnings
    Nov. 23
    • Cash
    • Common Stock
    • Paid-In Capital from Sale of Treasury Stock
    • Retained Earnings
    • Treasury Stock
    • Cash
    • Common Stock
    • Preferred Stock
    • Retained Earnings
    • Treasury Stock
    • Cash
    • Common Stock
    • Paid-In Capital from Sale of Treasury Stock
    • Preferred Stock
    • Retained Earnings

    Feedback

    b. What is the balance in Capital contributed to a corporation by the stockholders and others.Paid-In Capital from Sale of Treasury Stock on December 31 of the current year?
    $

    • Credit
    • Debit

    c. What is the balance in Treasury Stock on December 31 of the current year?
    $

    • Credit
    • Debit

    d. How will the balance in Treasury Stock be reported on the balance sheet?

    • As an asset
    • As an addition to stockholders' equity
    • As common stock in the stockholders' equity section
    • As a deduction from stockholders' equity
    • As a liability
Answer #1
a
Jan 31 Treasury Stock 255200 =8800*29
     Cash 255200
June. 14 Cash 170500 =5500*31
     Treasury Stock 159500 =5500*29
     Paid-In Capital from Sale of Treasury Stock 11000
Nov. 23 Cash 69300 =2100*33
     Treasury Stock 60900 =2100*29
     Paid-In Capital from Sale of Treasury Stock 8400
b Paid-In Capital from Sale of Treasury Stock 19400 Credit =11000+8400
c Treasury Stock 34800 Debit =255200-159500-60900
d As a deduction from stockholders' equity